Land,eToys is a profitable, medium-sized, retail company. Several years ago, it issued a 6.5% coupon bond, which pays interest semi-annually. The bond has a par value of $1,000 and will mature in ten years. It is currently priced in the market as $1,037.19 The average yields to maturity for 10-year corporate bonds are reported in the following table by bond rating: Yield (%) | Bond Rating Yield (%) 7.3 8.2 9.2 10.5 12.0 14.5 Bond Rating 5.4 5.7 6.5 Periodically, one company will purchase another by buying all of the target firms stock. The bonds of the target firm continue to exist. The debt obligation is assumed by the new firm. The credit risk of the bonds often changes because of this type of an event. Suppose that the firm Treasure Toys makes an announcement that they are purchasing LandoToys. Due to Treasure Toys projected financial structure after the purchase, Standard & Poors states that the bond rating for LandoToys bonds will change to BB. 4) Compute the current bond yield before and after the acquisition.
Answer #1
Current bond yield before acquisition = 65/1037.19 = 6.27%
Current bond yield after acquisition:
After acquisition the bond yield changes to 7.3%.
The bond price after acquisition = 1000/1.0365^20+32.50*(1.0365^20-1)/(0.0365*1.0365^20) = $     943.91
Current bond yield after acquisition = 65/943.91 = 6.89%